OpenStudy (anonymous):

? divied by 9=7 whats the mystery number?

OpenStudy (anonymous):

let's call ? x, ok? that would be x/9=7

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

then multiply by 9 on both sides to get x=(7*9)=63

OpenStudy (anonymous):

so, x=63=? which is your answer!

OpenStudy (anonymous):

ok thanks you:)

OpenStudy (anonymous):

you're quite welcome. can I have a medal please?

OpenStudy (anonymous):

how do i give a medal im new to this

OpenStudy (anonymous):

click next to one of my answers where it says "good answer"

OpenStudy (anonymous):

it dosent have that beside them

OpenStudy (anonymous):

over on the right side?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

maybe "give medal"? it's really ok...